# KAAE: Numerical Reasoning for Knowledge Graphs via Knowledge-aware Attributes Learning

### v1

@@ 3.2. Ordinal Knowledge Contrastive Learning
(10) $\bm{e}_{mix}^{t+}=\alpha\cdot\sum_{p\in Topk(\mathcal{E^{+}})}\bm{e}_{joint,p}^{att}+(1-\alpha)\cdot\bm{e}_{joint,h}^{att},$
(11) $\bm{e}_{mix}^{t-}=\beta\cdot\sum_{n\in Topk(\mathcal{E^{-}})}\bm{e}_{joint,n}^{att}+(1-\beta)\cdot\bm{e}_{joint,h}^{att},$
@@
(12) $\begin{split}\mathcal{L}_{CL}=-&\frac{1}{\mathcal{|G|}}\sum\nolimits_{l\in\mathcal{G}}\log\frac{exp(\textbf{p}^{\top}\textbf{u}/\tau)}{exp(\textbf{p}^{\top}\textbf{u}/\tau)+\sum\nolimits_{v_{-}\in\bm{e}_{mix}^{t-}}exp(\textbf{v}^{\top}\textbf{v}_{-}/\tau)},\end{split}$
where $\bm{p}=\bm{e}_{joint,h}^{att}+\bm{e}_{r}$, $\bm{u}=\bm{e}_{mix}^{t+}$, and $\tau$ is the temperature hyper-parameter. By implementing the strategy of randomly mixing the positive tails and employing head blending techniques (as described in Eqs. (10)-(11)), we notably increase the diversity of our samples. Consequently, this strategy significantly refines our model’s ability to navigate and interpret the nuances of semantic information, thereby reducing semantic ambiguity and bringing a more accurate and profound comprehension of semantics, advancing beyond the limitations imposed by the initial training data.
3.3. Training
@@
(13) $\begin{split}score(\bm{e}_{joint,h}^{att},\bm{e}_{r},\bm{e}_{joint,i}^{att})=&\varepsilon-\|\bm{e}_{joint,h}^{att}+\bm{e}_{r}-\bm{e}_{joint,i}^{att}\|_{1/2}+\\ &\Delta\|max(0,\bm{W}_{r}\bm{e}_{joint,h}-\bm{W}_{r}\bm{e}_{joint,i}^{att})\|^{2},\end{split}$
where $\Delta$ represents the weight in the numerical reasoning score function. The first term is derived from TransE (Eq. (17)), while the second term modifies the Order-embedding approach (Eq. (18)). $\bm{W}_{r}\in\mathbb{R}^{d_{emb}\times d_{emb}}$ is a projection matrix critical for mapping entities to a specific relational space, indicating that depending on the type of relation, each entity may assume a different order.
Loss Function. In this paper, we use the binary cross-entropy loss that presented in ConvE(Dettmers et al., 2018) and LiteralE(Kristiadi et al., 2019) so that we can convert the score of each triple obtained based on Eq.(13) to a probability by using a sigmoid function.Let $\mathcal{T}=\mathcal{G}\cup\mathcal{G}^{-}$ denote the training dataset,where $\mathcal{G}$ denotes the set of positive knowledge triples, $\mathcal{G}^{-}$ denotes the set of negative knowledge triples $\{(h,r,t^{{}^{\prime}})|h,t^{{}^{\prime}}\in\mathcal{E},r\in\mathcal{R},(h,r,t^{{}^{\prime}})\notin\mathcal{G}\}$. The binary-cross entropy loss is defined as follows:
(14) $\mathcal{L}_{BCE}=-\frac{1}{|\mathcal{T}|}\sum\limits_{l\in\mathcal{T}}(y_{l}\log(p_{l})+(1-y_{l})\log(1-p_{l})),$
@@
(15) $\mathcal{L}_{total}=\mathcal{L}_{BCE}+\lambda\mathcal{L}_{CL},$
@@
Table 1. ‘Numeric-aware’ indicates the availability of numeric information, while ‘Relation-aware’ indicates learning the presence of relation-specific. Knowledge-aware indicates Interacting with entities and relations simultaneously.
TransE BiGI R-GCN LiteralE RAKGE KAAE
[…]
Table 2. Results for numerical reasoning. Bold scores indicate the best results, while underlined scores represent the second-best results. The % of Improvement column shows the relative improvements of KAAE compared to the second-best scores.
@@
3.4. Theoretical Analysis
@@
To clarify how our model differs from these models, we compare our model with some popular models in Table1.
@@
Time Complexity Analysis. KAAE consists of three components, the MoEKA encoder, a process of contrastive learning, and the score function. In terms of MoEKA Encoder, the time complexity of processing every triple is $\mathcal{O}(|\mathcal{M}|)$. The time complexity of the numerical value embedding layer is $\mathcal{O}(|\mathcal{M}|)$ because we transform each attribute field into an embedding vector. In the joint perceptual attention layer, the query is the target relation and the number of keys and values corresponds to the number of attribute fields, so the time complexity remains $\mathcal{O}(|\mathcal{M}|)$. The gating layer of the MoEKA Encoder is not involved in the calculation of time complexity. During the process of contrastive learning, for every given triple(h,r,t), the time complexity is proportional to the number of positive samples $|\mathcal{P}[h,r]|$ and the number of negative samples $|\mathcal{N}[h,r]|$. The time complexity is $\mathcal{O}(|\mathcal{M}|\cdot rel\_{}num)$, where rel_num denotes the number of relations per entity. The TransE score function involves only addition and subtraction operations, so it does not change the time complexity. Therefore, the overall time complexity of KAAE is $\mathcal{O}(\mathcal{T}\cdot(|\mathcal{M}|+|\mathcal{M}|\cdot rel\_{}num))\approx\mathcal{O}(|\mathcal{T}|\cdot rel\_{}num)$, where $\mathcal{T}$ denotes the number of positive and negative triples used for training.
@@
The Expressiveness of Our Model. We conduct the expressiveness of numerical reasoning from the perspective of a one-hop reasoning task. Specifically, we aim to theoretically analyze the process of contrastive learning in KAAE and establish theoretical guarantees for the downstream performance of the learned representations. In the numerical reasoning, we denote the set of node representations as $\mathcal{V}$, and define the mean representations from one-hop neighborhoods of node $v$ as $\bm{z}=\frac{1}{\mathcal{N}(v)}\sum\nolimits_{\bm{u}\in\mathcal{N}(v)}\bm{u}$, where $\bm{u}$ represents the projected node representations of positive samples of node $v$, $\bm{z}$ describes the one-hop neighborhood pattern of node $v$. In the world, nodes belonging to the same semantic class tend to have similar neighborhood patterns, so $\bm{z}$ can be viewed sampled from $Z|Y\sim\mathcal{N}(\textbf{z}_{Y},I)$ where $Y$ is the latent semantic class indicating the one-hop pattern of node $v$. We demonstrate that minimizing KAAE’s objective in Equation (12) with an exponential moving average is equivalent to maximizing mutual information between representation $V$ and the one-hop pattern $Y$, which explains the rationality of KAAE in capturing one-hop patterns as follows:
(16) $\begin{split}\mathcal{L}_{CL}\geq H(V|Y)-H(V)=-I(V;Y),\end{split}$
@@
Eq. (16) indicates minimizing KAAE loss in Eq. (12) promotes maximizing the mutual information $I(V;Y)$ between representations and one-hop neighborhood context. In this way, our KAAE can exploit more latent semantics among KGs, hence facilitating numerical reasoning. All the detailed proofs are in Appendix B.
4. Experiments
4.1. Experimental Settings
@@
Datasets. We use three real-world KG datasets to evaluate the proposed model on a numerical reasoning task. Specifically, US-Cities [^1] https://simplemaps.com/data/us-cities is a large-scale city dataset containing basic information about several cities in the United States, including more than 70 attribute fields (e.g., household income, household education level, dimension,etc). Spotify is a knowledge graph provided by Spotify for developers [^2] https://www.kaggle.com/datasets/geomack/spotifyclassification. Credit is a knowledge graph constructed from default events in Taiwan(Yeh and Lien, 2009). Each credit card holder corresponds to an entity with numerical information about the credit card (e.g., total_delay, total_bill) as the numerical attribute. Additionally, it is worth noting that all three of the above datasets are modified to simulate numerical reasoning tasks under real conditions. Close to 20% of the numerical values are masked to zero (missing values) while using as few triples as possible for the training set. The remaining triples are used for evaluation. The statistics of the dataset can be found in Appendix C.
@@
Evaluation Metrics. We used the improved learning methods based on dropout, batch normalization, and linear transformation proposed alongside LTE(Zhang et al., 2022) for evaluation. Mean Reciprocal Rank (MRR), Mean Rank (MR) and Hits@{1, 3, 10} are reported as the performance metrics to evaluate the proposed model’s performance. Higher Hit@n and MRR values indicate a better performance, whereas lower MR values imply a better performance.
@@
Baseline. We compare KAAE with six groups of existing representation learning methods: Euclidean KGE including TransE(Bordes et al., 2013), ConvE(Dettmers et al., 2018), ComplEx (Trouillon et al., 2016), TuckER(Balazevic et al., 2019b), DistMult(Dong et al., 2014) and HAKE(Zhang et al., 2020). Hyperbolic KGE including MuRP(Balazevic et al., 2019a), ConE(Bai et al., 2021), AttH(Chami et al., 2020) and GIE(Cao et al., 2022b); GNNs-KGE including R-GCN(Schlichtkrull et al., 2018) and WGCN (Shang et al., 2019); Attributed KGE including KBLRN(García-Durán and Niepert, 2018), MT-KGNN (Tay et al., 2017) and LiteralE(Kristiadi et al., 2019); and Self-supervised model for graph including BiGI(Cao et al., 2021), SLiCE(Wang et al., 2021a), SimGCL(Yu et al., 2022), and RAKGE(Kim et al., 2023); Deterministic stand-alone value representation methods including DICE(Sundararaman et al., 2020), NEKG(Duan et al., 2021) and NRN(Bai et al., 2023). Among all the baselines, for those that do not take care of external numerical attributes, such as TransE, we follow their original approach and concat numerical values to the entity embeddings. The results of Self-supervised KGE are from RAKGE(Kim et al., 2023). More details for implements can refer to the Appendix C.
4.2. Experimental Results
Performance Comparison. To address these challenges, we introduce a new model termed KAAE (Knowledge-Aware Attributes Embedding), crafted for numerical reasoning via knowledge graphs. Specifically, to tackle semantic relevance, we utilize the Mixture-of-Experts (MoE) framework (Bian et al., 2023) to integrate semantic-guided knowledge experts, constructing expert networks for entities across various relational contexts. Additionally, we implement a joint interaction framework that leverages appropriate contexts and captures the intricate connections between entity-relation pairs and attributes within a unified vector space. Employing this knowledge-aware mechanism, our model discerns the significance and relevance of each attribute, effectively synthesizing the latent semantics among them. To address semantic ambiguity, diverging from prior work that primarily focuses on positive and negative samples, we introduce a novel ordinal knowledge contrastive learning method to generate high-quality ordinal samples. We optimize these ordinal samples based on cosine similarity, selecting the top-k samples with the highest similarity to enhance the learning of ordinal relations. This approach enables the model to capture the continuous features of numerical attributes and discern the subtle differences in boundary values, thus enhancing numerical reasoning. Furthermore, we provide theoretical guarantees for the effectiveness of our model. Extensive experiments substantiate the effectiveness of our model.
To validate the effectiveness of our ordinal knowledge contrastive learning method, we compared KAAE with those self-supervised models. MRR is improved from 0.331 for BiGI and 0.573 for RAKGE to 0.680 on Spotify dataset, we can deduce KAAE learn richer representations than graph-based approaches including BiGI, SLiCE, and SimGCL by considering extra numerical attributes of entities. Compared to RAKGE, KAAE focuses on generating and extracting hard high-quality ordinal samples to distinguish boundary numerical values and magnitudes accurately, which makes contrastive learning more effective in numerical reasoning tasks.
Besides, while HAKE, TuckER, and ConvE often outperform TransE and DistMutl models, Table 2 suggests otherwise. We speculate that this is because numerical reasoning is a one-to-one task in most cases, and is better suited to TransE and DistMult.
4.3. Experimental Analysis
Ablation Study. We conduct an ablation experiment to evaluate the performance of different variants of KAAE. Initially, we set $\delta_{e}$ in Eq.(3) to 0 to disable the entity-aware module. This modification allows the model to perceive numerical attributes solely through a relation-aware approach, without considering the embeddings of entities associated with the attributes. We also test the performance of KAAE without a full understanding of the entity semantics through the MoE module. Furthermore, we train the model without incorporating the top-k ordinal sampling methodology (Eqs. (10-11)) into the contrastive learning framework, which is essential for generating high-quality training samples and enhancing the model’s ability to discern discrete numerical boundaries. The results, as detailed in Table 3, indicate that the omission of any single module results in a decline in model performance. This finding emphasizes the critical role of making the most of entity-related information in numerical reasoning tasks and the significance of high-quality ordinal samples in contrastive learning.
Table 3. Ablation study of KAAE on Credit dataset. ‘MoE’ stands for Mixture-of-Experts, ‘KA’ stands for Knowledge-Aware, ‘OS’ stands for Ordinal-Sampling.
@@ 4.3. Experimental Analysis
[…]
Table 4. Case study on Credit dataset. IDXX means that we consider the information from an entity with a definite relation
Model Relation Entity Important Attributes(Attention Scores)
KAAE TOTAL_BILL_comp ID1763 TOTAL_BILL(0.278), LIMIT_BAL(0.243), AGE(0.193), TOTAL_DELAY(0.160), TOTAL_PAY(0.127)
MARRIAGE ID551 AGE(0.327), TOTAL_BILL(0.235), TOTAL_DELAY(0.207), TOTAL_PAY(0.144), LIMIT_BAL(0.088)
EDUCATION ID1130 AGE(0.356), TOTAL_DELAY(0.232), TOTAL_BILL(0.192), TOTAL_PAY(0.126), LIMIT_BAL(0.094)
RAKGE TOTAL_BILL_comp - AGE(0.203), TOTAL_PAY(0.202), TOTAL_BILL(0.201), LIMIT_BAL(0.197), TOTAL_DELAY(0.196)
MARRIAGE - TOTAL_PAY(0.202), TOTAL_BILL(0.2019), AGE(0.2017), TOTAL_DELAY(0.1971), LIMIT_BAL(0.1970)
EDUCATION - LIMIT_BAL(0.201), TOTAL_DELAY(0.200), TOTAL_PAY(0.1998), TOTAL_BILL(0.1997), AGE(0.1996)
Analysis of MoEKA Encoder. As mentioned above, the crucial role of our MoEKA Encoder is to extract the numerical information contained in the entity and relation with attributes, to accomplish joint perception with each attribute. The attention matrix, obtained using Eq. (5), is visualized in Figure 3. To make it easier to observe, we select a portion of all entities to visualize the attention scores, the first heatmap shows the entity-attribute association scores (setting $\delta_{r}$ in Eq. (3) to 0). The second heatmap shows the association scores after adding information from a relation to each entity (setting $\delta_{e}$ and $\delta_{r}$ to 0.5) in the first heatmap, effectively improving the problem of equalizing the attention to attributes for some entities. For example, Entity ID156 has close association scores with all 4 attributes except TOTAL_PAY, and taking into account the information of the relation TOTAL_PAY_comp, the MoEKA Encoder can better differentiate between these 5 attributes. In addition, we did another similar experiment to observe the effect of entity information on relation-attribute attention scores. The results are shown in Figure 7 in the Appendix, indicating that adding information from an entity to each relation effectively improves the problem of equalizing the attention to attributes in some relations.
@@
We also investigate the effect of the number of experts K in the MoE module, as depicted in Figure 4. It can be observed that the impact of the number of experts K on the final results generally follows a pattern of initial increase followed by a decrease, mainly affecting fine-grained metrics such as Hit@1 and MRR. Having either too many or too few experts is detrimental to the model’s learning performance. The model achieves the best results at K = 3.
Figure 3. Visualization of relevance scores of each numeric attribute without/with relation-aware on Credit dataset.
Figure 4. Research of Hyperparameter K on Spotify dataset.
Case Study. The MoEKA Encoder in KAAE was designed to obtain more robust attribute representations using the relevant context. Therefore, we investigated significant attributes contributed to the inference of a derived relation. Table 4 presents the attention scores of the MoEKA Encoder for all the attributes used to infer the comparison relations in the Credit dataset. When predicting relation TOTAL_BILL_comp, KAAE was able to focus more attention on the TOTAL_BILL attribute associated with TOTAL_BILL_comp. However, RAKGE(Kim et al., 2023) did not do so, paying essentially the same amount of attention to different attributes, which were all around 20%. Moreover, age is clearly a very important factor when predicting relations MARRIAGE and EDUCATION. By taking into account the information from entity embeddings, KAAE was able to focus more on the AGE attribute, while previous model RAKGE can not.

### v3

@@ Introduction
$\bm{e}_{mix}^{t+}=\alpha\cdot\sum_{p\in Topk(\mathcal{E^{+}})}\bm{e}_{joint,p}^{att}+(1-\alpha)\cdot\bm{e}_{joint,h}^{att}$ (10)
$\bm{e}_{mix}^{t-}=\beta\cdot\sum_{n\in Topk(\mathcal{E^{-}})}\bm{e}_{joint,n}^{att}+(1-\beta)\cdot\bm{e}_{joint,h}^{att}$ (11)
@@
$\begin{split}\mathcal{L}_{CL}=-&\frac{1}{\mathcal{|G|}}\sum\nolimits_{l\in\mathcal{G}}\log\frac{\varpi(\textbf{p}^{\top}\textbf{u}/\tau)}{\varpi(\textbf{p}^{\top}\textbf{u}/\tau)+\sum\nolimits_{v_{-}\in\bm{e}_{mix}^{t-}}\varpi(\textbf{v}^{\top}\textbf{v}_{-}/\tau)}\end{split}$ (13)
where $\bm{p}=\bm{e}_{joint,h}^{att}+\bm{e}_{r}$, $\bm{u}=\bm{e}_{mix}^{t+}$, $\tau$ is the temperature hyperparameter. $\varpi$ represents exponential function $exp(\cdot$). By implementing the strategy of randomly mixing the positive tails and employing head blending techniques (as described in Eqs. (10)-(11)), we notably increase the diversity of our samples and interpret the nuances of semantic information.
Training
@@
$\begin{split}&score(\bm{e}_{joint,h}^{att},\bm{e}_{r},\bm{e}_{joint,i}^{att})\\ =&\varepsilon-\|\bm{e}_{joint,h}^{att}+\bm{e}_{r}-\bm{e}_{joint,i}^{att}\|_{1/2}+\\ &\Delta\|max(0,\bm{W}_{r}\bm{e}_{joint,h}-\bm{W}_{r}\bm{e}_{joint,i}^{att})\|^{2}\end{split}$ (14)
where $\Delta$ represents the weight in the numerical reasoning score function. The first term is derived from TransE (DBLP:conf/nips/BordesUGWY13), while the second term modifies the Order-Embedding (DBLP:journals/corr/VendrovKFU15). $\bm{W}_{r}\in\mathbb{R}^{d_{emb}\times d_{emb}}$ is a projection matrix critical for mapping entities to a specific relational space, indicating that depending on the type of relation, each entity may assume a different order.
Loss Function. In this paper, we use the binary cross-entropy loss (DBLP:conf/aaai/DettmersMS018; DBLP:conf/semweb/KristiadiKL0F19) to optimize the training process. Let $\mathcal{T}=\mathcal{G}\cup\mathcal{G}^{-}$ denote the training dataset,where $\mathcal{G}$ denotes the set of positive knowledge triples, $\mathcal{G}^{-}$ denotes the set of negative knowledge triples $\{(h,r,t^{{}^{\prime}})|h,t^{{}^{\prime}}\in\mathcal{E},r\in\mathcal{R},(h,r,t^{{}^{\prime}})\notin\mathcal{G}\}$. The binary-cross entropy loss is defined as follows:
$\mathcal{L}_{BCE}=-\frac{1}{|\mathcal{T}|}\sum\limits_{l\in\mathcal{T}}(y_{l}\log(p_{l})+(1-y_{l})\log(1-p_{l}))$ (15)
@@
$\mathcal{L}_{total}=\mathcal{L}_{BCE}+\lambda\mathcal{L}_{CL}$ (16)
@@
Table 1: Results for numerical reasoning. Bold scores indicate the best results, while underlined scores represent the second-best results. The % of Improvement column shows the relative improvements of NumCoKE compared to the second-best scores.
Theoretical Analysis
@@
To clarify how NumCoKE differs from other models, we compare it with popular models in Table 13. (see Appendix).
@@
Time Complexity Analysis. NumCoKE consists of three components, the MoEKA encoder, a process of contrastive learning, and the score function. In terms of MoEKA Encoder, the time complexity of processing every triple is $\mathcal{O}(|\mathcal{M}|)$. The time complexity of the numerical value embedding layer is $\mathcal{O}(|\mathcal{M}|)$ because we transform each attribute field into an embedding vector. In the joint perceptual attention layer, the query is the target relation and the number of keys and values corresponds to the number of attribute fields, so the time complexity remains $\mathcal{O}(|\mathcal{M}|)$. The gating layer of the MoEKA Encoder is not involved in the calculation of time complexity. During the process of contrastive learning, for every given triple(h,r,t), the time complexity is proportional to the number of positive samples $|\mathcal{P}[h,r]|$ and the number of negative samples $|\mathcal{N}[h,r]|$. The time complexity is $\mathcal{O}(|\mathcal{M}|\cdot rel\_{}num)$, where rel_num denotes the number of relations per entity. The TransE score function involves only addition and subtraction operations, so it does not change the time complexity. Therefore, the overall time complexity of NumCoKE is $\mathcal{O}(\mathcal{T}\cdot(|\mathcal{M}|+|\mathcal{M}|\cdot rel\_{}num))\approx\mathcal{O}(|\mathcal{T}|\cdot rel\_{}num)$, where $\mathcal{T}$ denotes the number of positive and negative triples used for training.
@@
The Expressiveness of Our Model. We conduct the expressiveness of numerical reasoning from the perspective of a one-hop reasoning task. Specifically, we aim to theoretically analyze the process of contrastive learning in NumCoKE and establish theoretical guarantees for the downstream performance of the learned representations. In the numerical reasoning, we denote the set of node representations as $\mathcal{V}$, and define the mean representations from one-hop neighborhoods of node $v$ as $\bm{z}=\frac{1}{\mathcal{N}(v)}\sum\nolimits_{\bm{u}\in\mathcal{N}(v)}\bm{u}$, where $\bm{u}$ represents the projected node representations of positive samples of node $v$, $\bm{z}$ describes the one-hop neighborhood pattern of node $v$. In the world, nodes belonging to the same semantic class tend to have similar neighborhood patterns, so $\bm{z}$ can be viewed sampled from $Z|Y\sim\mathcal{N}(\textbf{z}_{Y},I)$ where $Y$ is the latent semantic class indicating the one-hop pattern of node $v$. We demonstrate that minimizing NumCoKE’s objective in Equation (13) with an exponential moving average is equivalent to maximizing mutual information between representation $V$ and the one-hop pattern $Y$, which explains the rationality of NumCoKE in capturing one-hop patterns as follows:
$\begin{split}\mathcal{L}_{CL}\geq H(V|Y)-H(V)=-I(V;Y)\end{split}$ (17)
@@
Eq. (17) indicates minimizing NumCoKE loss in Eq. (13) promotes maximizing the mutual information $I(V;Y)$ between representations and one-hop neighborhood context. In this way, our NumCoKE can exploit more latent semantics among KGs, hence facilitating numerical reasoning. All the detailed proofs are in Appendix B.
Experiment
Experimental Settings
@@
Datasets. We select three real-world KG datasets. Specifically, US-Cities [^1] https://simplemaps.com/data/us-cities contains basic information about cities in the United States. Spotify is a dataset of songs for developers [^2] https://www.kaggle.com/datasets/geomack. Credit (DBLP:journals/eswa/YehL09a) is a knowledge graph constructed from credit events in Taiwan. Additionally, all the three datasets are modified to simulate numerical reasoning tasks under real conditions. Close to 20% of the numerical values are masked to zero (missing values). More details and the statistics of datasets can be found in Appendix C.
@@
Evaluation Metrics. We used the improved learning methods based on dropout, batch normalization, and linear transformation proposed alongside LTE (DBLP:conf/www/Zhang0YW22) for evaluation. Mean Reciprocal Rank (MRR), Mean Rank (MR) and Hits@{1, 3, 10} are reported as metrics. Higher Hit@n and MRR values indicate a better performance, whereas lower MR values imply a better performance.
@@
Baseline. We consider six groups of 23 baseline methods: Euclidean KGE including TransE (DBLP:conf/nips/BordesUGWY13), ConvE (DBLP:conf/aaai/DettmersMS018), TuckER (DBLP:conf/emnlp/BalazevicAH19), HAKE (DBLP:conf/aaai/ZhangCZW20), DaBR (DBLP:conf/coling/WangLBG25), KGDM (DBLP:conf/aaai/LongZLWLW24), and RuIE (DBLP:conf/coling/LiaoDH025). Hyperbolic KGE including MuRP (DBLP:conf/nips/BalazevicAH19), ConE (DBLP:conf/nips/BaiYRL21), AttH (DBLP:conf/acl/ChamiWJSRR20), GIE (DBLP:conf/aaai/CaoX0CH22) and LorentzKG (DBLP:conf/acl/Fan0CCDY24); GNNs-KGE including R-GCN (DBLP:conf/esws/SchlichtkrullKB18) and WGCN (DBLP:conf/aaai/ShangTHBHZ19); Attributed KGE including KBLRN (DBLP:conf/uai/Garcia-DuranN18), MT-KGNN (DBLP:conf/cikm/TayTPH17) and LiteralE (DBLP:conf/semweb/KristiadiKL0F19); Self-supervised model for graph including BiGI (DBLP:conf/wsdm/CaoLGLL021), SLiCE (DBLP:conf/www/WangAHCR21), SimGCL (DBLP:conf/sigir/YuY00CN22), and RAKGE (DBLP:conf/kdd/KimKKPJP23); Deterministic value representation methods including NEKG (DBLP:conf/emnlp/DuanYT21) and NRN (DBLP:conf/kdd/BaiLLYYS23). Among all the baselines, for those that do not take care of external numerical attributes, such as TransE, we follow the original approach and concat the embeddings of entity and attributes. The results of self-supervised KGE are from RAKGE. More details can refer to the Appendix C.
Experimental Results
The main results are detailed in Table1. We conduct 5 rounds of experiments and take the average as the final result, which reveals that NumCoKE makes significant progress in all the metrics and achieves new SOTA results. All the evaluated models are implemented on a server with one GPU (NVIDIA A800, 80GB). Notably, MRR on Credit increased from 0.493 for TransE and 0.498 for DaBR to 0.794, which highlights NumCoKE’s robust capability in capturing the asymmetric nature of relations and the ordinal information. At the same time, NumCoKE’s MRR significantly outperforms to hierarchy-aware models such as TuckER and hyperbolic KGE methods, affirming its superior utilization of the hierarchical structures inherent in numerical reasoning relations. RuIE performs weakly, which we infer is due to excessive focus on logical rules rather than comparative relationships between numerical semantics. Although utilizing the extra attributes of entities, attribute-based models like KBLRN and MT-KGNN experience performance declines due to the relevance of numerical attributes, whereas NumCoKE excels under these conditions. Deterministic approaches such as NEKG and NRN are unable to recognize semantic differences between numerical fields and only consider the distribution of attribute values in each field (e.g., 51kg and 51 years old), whereas NumCoKE sets up different representation for each field and utilizes the knowledge perceptual attention mechanism that allows it to focus on attributes that are more relevant to the relation. In addition, compared to LiteralE and RAKGE, NumCoKE excels by fully exploiting the potential of the relational contexts and the interactions among entities, relations and numerical attributes. More details can refer to the study for ordinal knowledge contrastive learning in Appendix D.
Experimental Analysis
Ablation Study. To assess the contribution of each component in NumCoKE, we conduct ablation studies by removing key modules individually. (1) Knowledge-aware attention (KA): We disable the KA module by setting $\delta_{e}=\delta_{r}=0$ in Eq. (3), effectively removing the influence of relation- and entity-specific signals on attribute modeling. As shown in Table 2, this leads to a significant performance drop (↓8.1% on Credit), confirming the importance of incorporating both entity and relation semantics for numerical reasoning. (2) Mixture-of-experts (MoE): We remove the MoE mechanism and instead use static entity representations across all relation contexts. This results in a 5.0% accuracy drop, indicating that dynamically adapting entity embeddings based on relational context is crucial for capturing semantic nuances. (3) Ordinal sampling (OS): We further ablate the top-$k$ ordinal sampling strategy used in Eq. (10)–(11). Without OS, the model relies solely on randomly sampled contrastive pairs, resulting in degraded performance (↓6.4% on Credit). This highlights the effectiveness of OS in improving the model’s ability to distinguish subtle ordinal differences among close numerical values. We also conduct a single-module-addition study, the results can be seen in the Appendix.
@@
$\times$ $\surd$ $\surd$ 0.712 0.764 0.851 9 0.765
$\surd$ $\times$ $\surd$ 0.688 0.763 0.846 10 0.740
$\surd$ $\surd$ $\times$ 0.698 0.776 0.843 10 0.750
$\surd$ $\surd$ $\surd$ 0.745 0.818 0.888 7 0.794
Table 2: Ablation study on Credit dataset.
@@
Analysis of MoEKA Encoder. We investigate the effect of the number of experts K in the MoE module, as depicted in Figure 3. It can be observed that the impact of the number of experts K on the final results generally follows a pattern of initial increase followed by a decrease, mainly affecting fine-grained metrics such as Hit@1 and MRR. Having either too many or too few experts is detrimental to the model’s learning performance, this is because the same relational category corresponds to similar contexts, for example, ranking_comp and disabled_comp are both comparative relations, and too many experts can cause semantic confusion.
Figure 3: Research of Hyperparameter K on Spotify dataset.
Effect of Relation-Entity Perception Weights. To assess the relative contributions of entities and relations in numerical attribute modeling, we vary their attention weights in the joint perception mechanism (Eq. 3). Specifically, we adjust the proportion assigned to relations from 0% to 100% and evaluate the performance of the model on the Credit dataset. As shown in Figure 4, NumCoKE achieves optimal performance when relations contribute 40%–90% of the total attention weight, with peak results at 90%. Compared to the entity-only variant (0% relation weight), this configuration improves MRR by 46.1% (from 0.186 to 0.272), demonstrating the significant value of relation semantics in attribute perception. However, performance deteriorates when relations are used exclusively (i.e., 100%), mirroring the trend observed in RAKGE (DBLP:conf/kdd/KimKKPJP23), which neglects entity semantics. These results highlight the necessity of jointly modeling entity and relation semantics to effectively capture numerical attributes.